OpenStudy (jack1):

if i have a circle of radius 1 what are the dimensions for a square with the same area as the circle?

OpenStudy (jack1):



OpenStudy (anonymous):

Side length: \(\sqrt{\text{Area}}\)

OpenStudy (anonymous):

So the square would be \(\sqrt{\pi} \times \sqrt{\pi}\)

OpenStudy (jack1):

cool, cheers, so area = pi then?

OpenStudy (anonymous):

Yes \(\pi r^2\) and the radius is 1 for the area of the circle.

OpenStudy (jack1):

sweet, so this holds true for circles of all radius sizes? (ie side legnth = sqrt pi x r)?

OpenStudy (anonymous):